OpenStudy (anonymous):

simplify (40^3+4a^2+33a+27)/5a+3

OpenStudy (therealmeeeee):

lets see wat i can do

OpenStudy (therealmeeeee):

5a x 4a = 20a 2/3 + 3/2 = 13/12 Answer : 20a^13/12

OpenStudy (therealmeeeee):

do u get that

OpenStudy (anonymous):

the first term should be 40a^3

OpenStudy (therealmeeeee):

yes

OpenStudy (therealmeeeee):

did u get it

OpenStudy (anonymous):

no

OpenStudy (therealmeeeee):

were r u stuck

OpenStudy (therealmeeeee):

Answer : 20a^13/12

OpenStudy (therealmeeeee):

now do u get it

OpenStudy (therealmeeeee):

medal me

OpenStudy (anonymous):

You could use polynomial long division. It's not hard to do. It's a lot like normal long division.

OpenStudy (anonymous):

It's like if 20/4 = 5. The result of using polynomial long division turns the question into 8a^2 -4a+9.

OpenStudy (anonymous):

It divides out quite nicely. There turns out to be no remainder. I'm pretty sure that they want you to simplify it using polynomial long division.